OpenStudy (anonymous):

2. 7m − 5 − 13m = 25

OpenStudy (anonymous):

m= -5

OpenStudy (anonymous):

show work pls

OpenStudy (anonymous):

first add the like terms

OpenStudy (anonymous):

2.7m - 5 - 13m = 25 -10.3m - 5 = 25 add ten to both sides -10.3m = 30 divide by -10.3 on both sides m = -2.91262136

OpenStudy (anonymous):

7m−5−13m=25 add like terms so it becomes -5-6m=25 Add 5 to both -5 and 25 the -5 cancels the 25 becomes 30 30 divided by -6 = -5

OpenStudy (anonymous):